# Radiative corrections to superallowed $\boldsymbol{\beta}$ decays in effective field theory

### v1

@@ Abstract
$|V_{ud}|^{2}+|V_{us}|^{2}+|V_{ub}|^{2}=1,$ (1)
which, at the moment, displays a tension at the level of $2.8\sigma$ Cirigliano et al. (2023a). While a separate tension among $V_{us}$ determinations from kaon decays could potentially be resolved by future measurements at NA62 Cirigliano et al. (2023a), also $V_{ud}$ has come under increased scrutiny in recent years, mainly in view of the increased tension that followed from a reevaluation of universal radiative corrections (RC) associated with $\gamma W$ box diagrams Marciano and Sirlin (2006); Seng et al. (2018, 2019); Czarnecki et al. (2019); Seng et al. (2020); Hayen (2021); Shiells et al. (2021). Such a violation of CKM universality could point to a wide range of possible beyond-the-Standard-Model scenarios Belfatto et al. (2020); Coutinho et al. (2020), including vector-like quarks Cheung et al. (2020); Belfatto and Berezhiani (2021); Branco et al. (2021) and leptons Crivellin et al. (2020); Kirk (2021), could be interpreted as a modification of the Fermi constant Marciano (1999); Crivellin et al. (2021a), the violation of lepton flavor universality Crivellin and Hoferichter (2020); Crivellin et al. (2021b); Capdevila et al. (2021); Crivellin and Hoferichter (2021); Crivellin et al. (2021c); Marzocca and Trifinopoulos (2021), or, more generally, in the context of Standard-Model EFT Alok et al. (2023); Cirigliano et al. (2022a); Cirigliano et al. (2024); Dawid et al. (2024). It is thus paramount to consolidate the evaluation of $V_{ud}$ and potentially even improve its precision.
@@
$f$ is a phase-space factor that includes the Fermi function, due to the Coulomb interaction of the outgoing electron in the nuclear field, the nuclear electroweak (EW) form factor, nuclear recoil, atomic electron screening, and atomic overlap Hardy and Towner (2020); Gorchtein and Seng (2024). The other terms denote purely theoretical input due to isospin-breaking and non-Coulomb RC. $\delta_{C}$ denotes the deviation of the Fermi matrix element, $M_{\text{F}}=\langle f|\tau^{+}|i\rangle=M_{\text{F}}^{(0)}(1-\delta_{C}/2)$ from its isospin-limit value $M_{\text{F}}^{(0)}=\sqrt{2}$. The so-called outer correction $\delta_{R}^{\prime}$ encodes all infrared-sensitive RC not included in the Fermi function. The remaining RC are collectively denoted as the inner correction and are usually split into the single-nucleon correction $\Delta^{V}_{R}$ and the nuclear-structure-dependent term $\delta_{\text{NS}}$, which arises from the phase-space average of a correction that in general depends on the positron energy $E_{e}$, as pointed out recently in Refs. Gorchtein (2019); Seng and Gorchtein (2023a).
Currently, the largest uncertainties reside in $\delta_{C}$ and $\delta_{\text{NS}}$. First, control over $\delta_{C}$ has long been a concern Miller and Schwenk (2008, 2009) and Refs. Martin et al. (2021); Condren and Miller (2022); Seng and Gorchtein (2023b); Crawford and Miller (2022); Seng and Gorchtein (2024a, b) provide recent studies and strategies for improvements. Second, while the single-nucleon, universal RC from $\gamma W$ box diagrams have reached a good level of maturity, including a comprehensive analysis in EFT Cirigliano et al. (2023b) and a first lattice-QCD evaluation Ma et al. (2024), the same cannot be said for the nucleus-dependent effects of the same diagrams, called $\delta_{\text{NS}}$, whose estimate dominates the resulting uncertainty in $V_{ud}$ Gorchtein (2019). The formalism for an evaluation using dispersion relations has been put forward in Refs. Seng and Gorchtein (2023a); Gorchtein and Seng (2024), including subtleties that arise in the case of low-lying intermediate states, such as the $3^{+}$ and $1^{+}$ levels of 10B in the ${}^{10}\text{C}\to{}^{10}\text{B}$ $0^{+}\to 0^{+}$ transition.
@@
to systematically expand the $\beta$ decay amplitude in the ratios of scales probed by the virtual photon,
@@
besides the ratio $G_{F}q^{2}_{\text{ext}}$ that sets the overall scale. Our goal is to catalog all corrections to superallowed $\beta$ decays up to order $\mathcal{O}(\alpha\epsilon_{\chi})$ and $\mathcal{O}(\alpha\epsilon_{\not{\pi}})$, which are needed for extracting $V_{ud}$ at the $10^{-4}$ level.
@@
In terms of the nucleon $N^{T}=(p,n)$ isodoublet, the nucleon four-velocity $v_{\mu}$ and spin $S_{\mu}$, and isospin Pauli matrices $\tau^{a}$, the leading-order (LO) EW one-body (1b) Lagrangian is
@@
where the ellipsis denotes omitted terms involving pion fields or of higher order in $\epsilon_{\chi}$. The effects of hard photons with virtuality $Q^{2}\geq\Lambda_{\chi}^{2}$ are captured in the deviation of the vector coupling $g_{V}$ from one (and $g_{A}$ from $g_{A}^{\text{QCD}}$ Cirigliano et al. (2022b)), see Ref. Cirigliano et al. for explicit expressions. Hard photons also generate EW two-body (2b) contact operators. We can write two $S$-wave operators relevant for superallowed $\beta$-decays that connect ${}^{1}S_{0}$ to ${}^{1}S_{0}$ states, with isospin $I=1$ and $I=2$, given by
@@
Weinberg power counting based on naive dimensional analysis would indicate that $g^{N\!N}_{V1,V2}=\mathcal{O}(\Lambda_{\chi}^{-3})$, but the requirement of regulator independence promotes the LECs to $\mathcal{O}(\Lambda_{\chi}^{-1}F_{\pi}^{-2})$, where $F_{\pi}=92.3\,\text{MeV}$ is the pion decay constant. The values of $g^{N\!N}_{V1,V2}$ are not known, but we will discuss strategies to obtain them below.
Figure 1: Representative diagrams for RC to superallowed $\beta$ decays up to $\mathcal{O}(\alpha\epsilon_{\chi})$ and $\mathcal{O}(\alpha\epsilon_{\not{\pi}})$. Leptons, nucleons, photons, and pions are denoted by plain, double, wavy, and dashed lines, respectively. A blue circle denotes the insertion of the EW current, including $\mathcal{O}(\alpha)$ corrections from hard photon exchange. Black circles denotes 1b EW and EM currents. The red and green ovals denote the wave functions of the initial and final nuclei, the blue oval represents the iteration of the nuclear interaction.
Within this chiral EFT with dynamical photons and leptons we compute EW transition amplitudes involving multiple nucleons, see Fig. 1 for some of the topologies relevant for nuclear decays. In the presence of more than one nucleon, the photon four-momentum can be in three regions:
@@
In the nuclear EFT, potential and soft modes are integrated out and give rise to an EW transition operator, analogous to the pion-exchange potential in the strong sector. In addition, hard photons also contribute short-range transition operators proportional to $g^{N\!N}_{V1,V2}$ in Eq. (6), see diagram 1$(e)$. Because the exchange of a hard, soft, or potential photon leaves the intermediate nuclear state far off-shell, these contributions can be calculated by taking the matrix element of the transition operator between the wave functions of the initial and final state. On the other hand, ultrasoft photons are sensitive to nuclear excitations and to the spectrum of intermediate states that are connected to the initial and final state by EW and EM currents.
@@
Ultrasoft modes: Ultrasoft modes contribute at $\mathcal{O}(\alpha)$ through the LO photon–nucleon coupling. Through topologies such as 1$(b)$ and 1$(c)$ (and real emission topologies that we have omitted), ultrasoft modes give rise to the Sirlin function Sirlin (1967) and reproduce the $\mathcal{O}(\alpha)$ expansion of the Fermi Fermi (1934) function, with the correct nuclear charge $Z$. Subleading interactions, such as the interactions of the photon with the nucleon magnetic moment, are proportional to the ultrasoft momentum and appear at $\mathcal{O}(\alpha\epsilon_{\text{recoil}})$ beyond the order at which we work.
Potential modes: Through topology 1$(c)$, potential modes give rise to $\mathcal{O}(\alpha\epsilon_{\not{\pi}})$ and $\mathcal{O}(\alpha\epsilon_{\chi})$ corrections to $\delta_{\text{NS}}$. The former depend on the electron energy ($E_{e}$) and are induced by diagrams with the EW vector current and the EM charge density. The latter are $E_{e}$-independent and are induced by the axial current and the nucleon magnetic moments or recoil corrections to the vector current. Three-body (3b) potential contribute at $\mathcal{O}(\alpha\epsilon_{\chi}^{2})$ and are not shown.
@@
Hard modes: Hard modes give $\mathcal{O}(\alpha)$ corrections to $g_{V}$ Cirigliano et al. (2023b) and generate the $\mathcal{O}(\alpha\epsilon_{\chi})$ two-nucleon counterterms ($g^{N\!N}_{V1,V2}$) needed for renormalization. In addition, they produce $\mathcal{O}(\alpha\epsilon_{\not{\pi}},\alpha\epsilon_{\chi})$ effects in $\delta_{\text{NS}}$ through the electromagnetic pion mass splitting in pion-mediated 2b currents, see diagram 1$(d)$. These corrections are the nuclear analogs of the pion-induced RC in neutron decay Cirigliano et al. (2022b).
The implication of this analysis is that in chiral EFT the dominant contribution to $\delta_{\text{NS}}$ comes from the matrix element of appropriate EW potentials between the initial and final nuclear states, where some contributions (from pion exchange) do not arise from nuclear $\gamma W$ box diagrams. Sensitivity of $\delta_{\text{NS}}$ to intermediate nuclear states, including low-lying levels, arises from ultrasoft contributions that start to $\mathcal{O}(\alpha\epsilon_{\text{recoil}})$. This result is seemingly at odds with a recent dispersive analysis Seng and Gorchtein (2023a); Gorchtein and Seng (2024) in which some individual contributions scale as $\mathcal{O}(\alpha\sqrt{\epsilon_{\text{recoil}}})$, and thus enhanced compared to the identified EFT scalings. We therefore turn next to a detailed comparison to the dispersive representation.
@@
One way to perform the loop integral in Eq. (Radiative corrections to superallowed $\boldsymbol{\beta}$ decays in effective field theory) relies on a Wick rotation $\nu\to i\nu_{E}$, which is advantageous when the Compton tensor is expressed via a dispersion relation. However, as pointed out in Ref. Gorchtein and Seng (2024), such a Wick rotation is not always possible, because in the case of low-lying nuclear states, as does happen in the ${}^{10}\text{C}\to{}^{10}\text{B}$ decay, the additional pole can move into the third quadrant and thus must be subtracted explicitly. It was found that this contribution becomes singular for $E_{e}\to 0$, which could lead to a numerical enhancement. Such an enhancement should be reflected by the momentum scaling and a different region in the EFT analysis.
@@
where $s=M^{2}+\nu^{2}-\mathbf{q}^{2}+2M\nu$ and $M^{2}-\bar{M}^{2}=2M\Delta$. Here, $g_{A}$ and $g_{M}$ parameterize the matrix elements for the interaction with the EW and EM current, respectively. We focus on a single intermediate state with mass $\bar{M}$, with $\Delta>0$ corresponding to a low-lying state. The prefactor has been chosen to match the corresponding EFT expression Cirigliano et al. , counting the binding energy $\Delta\simeq q_{\text{ext}}$ as before.
We can evaluate the integral by collecting all three residues in the upper half plane, see Ref. Cirigliano et al. , which gives
@@
which is again finite for $E_{e}\to 0$, but, contrary to Eq. (12), scales as $\mathcal{O}(\alpha\sqrt{\epsilon_{\text{recoil}}})$ and could thus be enhanced numerically. The solution to this apparent mismatch is that the Wick-rotated integral also involves terms scaling with $\sqrt{\Delta}$, and one can show explicitly that Cirigliano et al.
@@
Leading contributions to $\delta_{\emph{NS}}$.—In Ref. Cirigliano et al. we derive the nuclear decay rate in the EFT framework, while here we focus on the implications for $\delta_{\text{NS}}$. Potential modes induce an effective Hamiltonian of the form
@@
with the explicit expressions for $\mathcal{V}_{E}$, $\mathcal{V}_{E}^{\pi}$, and $\mathcal{V}_{m_{e}}^{\pi}$ given in Ref. Cirigliano et al. . The latter two depend on the pion mass splitting, $M_{\pi^{\pm}}^{2}-M_{\pi^{0}}^{2}=2e^{2}F_{\pi}^{2}Z_{\pi}$, encoding effects of hard photons. These potentials are energy dependent and affect both the spectral shape and the total decay rate at $\mathcal{O}(\alpha\epsilon_{\not{\pi}})$.
Instead of inserting a lepton momentum, we can consider subleading vertices in diagrams 1$(c,d)$. One order down in the chiral expansion at $\mathcal{O}(\alpha\epsilon_{\chi})$ we obtain potentials that are independent of the lepton momenta and thus contribute to $\mathcal{V}^{0}$. These $\mathcal{O}(\alpha\epsilon_{\chi})$ terms can be further decomposed as
@@
corresponding to diagrams 1$(c,d)$ via magnetic, recoil, and contact-term contributions Cirigliano et al. . Beyond tree level, the potentials (16) and (17) receive corrections from soft pions and photons at $\mathcal{O}(\alpha\epsilon_{\chi}^{2})$ and $\mathcal{O}(\alpha^{2})$, beyond the accuracy of this work. At this order we also expect effects from 3b potentials.
@@
$\mathcal{V}_{0}^{\text{CT}}$ depends on two unknown LECs and corresponds to genuine new 2b contributions arising from high-momentum photon exchange. It is an intrinsic two-nucleon effect that cannot be obtained from one-nucleon processes.
In the EFT approach, up to the order considered, $\delta_{\text{NS}}$ is entirely determined by matrix elements of appropriate potentials, see Eq. (15), between the initial and final states without dependence on intermediate nuclear states. The EFT power counting indicates that $\delta_{\text{NS}}$ receives a LO $E_{e}$-independent contribution of $\mathcal{O}(\alpha\epsilon_{\chi})$, $\delta_{\text{NS}}^{(0)}$, and an $E_{e}$-dependent contribution of $\mathcal{O}(\alpha\epsilon_{\not{\pi}})$, $\delta^{E}_{\text{NS}}$. In the case of $\delta_{\text{NS}}^{(0)}$ we also found an $\mathcal{O}(\alpha^{2})$ potential $\mathcal{V}_{+}$ that needs to be included for $\mathcal{O}(10^{-4})$ precision Cirigliano et al. .

### v2

@@ Abstract
$|V_{ud}|^{2}+|V_{us}|^{2}+|V_{ub}|^{2}=1.$ (1)
At the moment, Eq. (1) displays a tension at the level of $2.8\sigma$ Cirigliano et al. (2023a). While a separate tension among $V_{us}$ determinations from kaon decays could potentially be resolved by future measurements at NA62 Cirigliano et al. (2023a), also $V_{ud}$ has come under increased scrutiny in recent years, mainly in view of the increased tension that followed from a reevaluation of universal radiative corrections (RC) associated with $\gamma W$ box diagrams Marciano and Sirlin (2006); Seng et al. (2018, 2019); Czarnecki et al. (2019); Seng et al. (2020); Hayen (2021); Shiells et al. (2021). Such a violation of CKM unitarity could point to a wide range of possible beyond-the-Standard-Model scenarios Belfatto et al. (2020); Coutinho et al. (2020), including vector-like quarks Cheung et al. (2020); Belfatto and Berezhiani (2021); Branco et al. (2021); Crivellin et al. (2021a) and leptons Crivellin et al. (2020); Kirk (2021), could be interpreted as a modification of the Fermi constant Marciano (1999); Crivellin et al. (2021b), the violation of lepton flavor universality Crivellin and Hoferichter (2020); Crivellin et al. (2021c); Capdevila et al. (2021); Crivellin and Hoferichter (2021); Crivellin et al. (2021d); Marzocca and Trifinopoulos (2021), or, more generally, in the context of Standard-Model EFT Alok et al. (2023); Cirigliano et al. (2022a, 2024a); Dawid et al. (2024). It is thus paramount to consolidate the evaluation of $V_{ud}$ and potentially even improve its precision.
@@
$f$ is a phase-space factor that includes the Fermi function, due to the Coulomb interaction of the outgoing electron in the nuclear field, the nuclear electroweak (EW) form factor, nuclear recoil, atomic electron screening, and atomic overlap Hardy and Towner (2020); Gorchtein and Seng (2024). The other terms denote purely theoretical input due to isospin-breaking and non-Coulomb RC. $\delta_{C}$ denotes the deviation of the Fermi matrix element, $M_{\text{F}}=\langle f|\tau^{+}|i\rangle=M_{\text{F}}^{(0)}(1-\delta_{C}/2)$ from its isospin-limit value $M_{\text{F}}^{(0)}=\sqrt{2}$. The so-called outer correction $\delta_{R}^{\prime}$ encodes all infrared-sensitive RC not included in the Fermi function. At ${\mathcal{O}}(\alpha)$, these include the Sirlin function Sirlin (1967). The precise extraction of $V_{ud}$ requires control of corrections of ${\mathcal{O}}(\alpha^{2}Z)$ and higher Jaus (1972); Jaus and Rasche (1987); Sirlin and Zucchini (1986). The remaining RC are collectively denoted as the inner correction and are usually split into the single-nucleon correction $\Delta^{V}_{R}$ and the nuclear-structure-dependent term $\delta_{\text{NS}}$. The latter arises from the phase-space average of a correction that in general depends on the positron energy $E_{e}$, as pointed out recently in Refs. Gorchtein (2019); Seng and Gorchtein (2023a).
Currently, the largest uncertainties reside in $\delta_{C}$ and $\delta_{\text{NS}}$. First, control over $\delta_{C}$ has long been a concern Miller and Schwenk (2008, 2009) and Refs. Martin et al. (2021); Condren and Miller (2022); Seng and Gorchtein (2023b); Crawford and Miller (2022); Seng and Gorchtein (2024a, b) provide recent studies and strategies for improvements. Second, while the single-nucleon, universal RC from $\gamma W$ box diagrams have reached a good level of maturity, including a comprehensive analysis in EFT Cirigliano et al. (2023b) and a first lattice-QCD evaluation Ma et al. (2024), the same cannot be said for the nucleus-dependent effects of the same diagrams. The nuclear correction called $\delta_{\text{NS}}$ dominates the resulting uncertainty in $V_{ud}$ Gorchtein (2019). The formalism for an evaluation using dispersion relations has been put forward in Refs. Seng and Gorchtein (2023a); Gorchtein and Seng (2024), including subtleties that arise in the case of low-lying intermediate states, such as the $3^{+}$ and $1^{+}$ levels of 10B in the ${}^{10}\text{C}\to{}^{10}\text{B}$ $0^{+}\to 0^{+}$ transition Gennari et al. (2024).
@@
to systematically expand the $\beta$ decay amplitude in the ratios of scales probed by the virtual photon. Besides the ratio $G_{F}q^{2}_{\text{ext}}$ that sets the overall scale, these are
@@
scaling roughly as $\simeq 0.005$, $\simeq 0.05$, and $\simeq 0.1$, respectively. Our goal is to catalog all corrections to superallowed $\beta$ decays at the permille level. This requires keeping $\mathcal{O}(\alpha\epsilon_{\chi})$ and $\mathcal{O}(\alpha\epsilon_{\not{\pi}})$ corrections, which are the focus of this Letter. Terms that are subleading in $\alpha$, but enhanced by the nuclear charge $Z$ or large logarithms, e.g., $\mathcal{O}(Z\alpha^{2})$ or $\mathcal{O}(\alpha^{2}\log r)$, with $r$ a ratio of the scales in Eq. (3), are also relevant and discussed in detail in Ref. Cirigliano et al. (2024b), as is the potential role of $\mathcal{O}(\alpha\epsilon_{\chi}^{2})$ corrections, which are not yet included at present.
@@
In terms of the heavy-baryon nucleon $N^{T}=(p,n)$ isodoublet, the nucleon four-velocity $v_{\mu}$ and spin $S_{\mu}$, and isospin Pauli matrices $\tau^{a}$ Jenkins and Manohar (1991); Bernard et al. (1995), the leading-order (LO) EW one-body (1b) Lagrangian is
@@
where the ellipsis denotes omitted terms involving pion fields or of higher order in $\epsilon_{\chi}$. The effects of hard photons with virtuality $Q^{2}\geq\Lambda_{\chi}^{2}$ are captured in the deviation of the vector coupling $g_{V}$ from one (and $g_{A}$ from $g_{A}^{\text{QCD}}$ Cirigliano et al. (2022b)), see Ref. Cirigliano et al. (2024b) for explicit expressions. Hard photons also generate EW two-body (2b) contact operators. We can write two $S$-wave operators relevant for superallowed $\beta$-decays that connect ${}^{1}S_{0}$ to ${}^{1}S_{0}$ states, with isospin $I=1$ and $I=2$, given by
@@
Weinberg power counting based on naive dimensional analysis would indicate that $g^{N\!N}_{V1,V2}=\mathcal{O}(\Lambda_{\chi}^{-3})$, but the requirement that the final nuclear amplitude be independent of the regulator promotes the low-energy constants (LECs) to $\mathcal{O}(\Lambda_{\chi}^{-1}F_{\pi}^{-2})$, where $F_{\pi}=92.3\,\text{MeV}$ is the pion decay constant. The values of $g^{N\!N}_{V1,V2}$ are not known, but we will discuss strategies to obtain them below.
Figure 1: Representative diagrams for RC to superallowed $\beta$ decays up to $\mathcal{O}(\alpha\epsilon_{\chi})$ and $\mathcal{O}(\alpha\epsilon_{\not{\pi}})$. Leptons, nucleons, photons, and pions are denoted by plain, double, wavy, and dashed lines, respectively. A blue circle denotes the insertion of the EW current, including $\mathcal{O}(\alpha)$ corrections from hard photon exchange. Black circles denote 1b EW and EM currents. The red and green ovals denote the wave functions of the initial and final nuclei, the blue oval represents the iteration of the nuclear interaction.
Within this chiral EFT with dynamical photons and leptons we compute EW transition amplitudes involving multiple nucleons, see Fig. 1 for some of the topologies relevant for nuclear decays. In the presence of more than one nucleon, the photon four-momentum can be in three regions, see, e.g., Refs. Beneke and Smirnov (1998); Smirnov (2002):
@@
In the nuclear EFT, potential and soft modes are integrated out and give rise to an EW transition operator, analogous to the pion-exchange potential in the strong sector. In addition, hard photons also contribute to short-range transition operators proportional to $g^{N\!N}_{V1,V2}$ in Eq. (6), see diagram 1$(e)$. Because the exchange of a hard, soft, or potential photon leaves the intermediate nuclear state far off-shell, these contributions can be calculated by taking the matrix element of the transition operator between the wave functions of the initial and final state. On the other hand, ultrasoft photons are sensitive to nuclear excitations and to the spectrum of intermediate states that are connected to the initial and final state by EW and EM currents.
@@
Ultrasoft modes: Ultrasoft modes contribute at $\mathcal{O}(\alpha)$ through the LO photon–nucleon coupling. Through topologies such as 1$(b)$ and 1$(c)$ (and real emission topologies that we have omitted), ultrasoft modes give rise to the Sirlin function Sirlin (1967) and reproduce the $\mathcal{O}(\alpha)$ expansion of the Fermi Fermi (1934) function, with the correct nuclear charge $Z$, see Ref. Cirigliano et al. (2024b) for details. Using known results, terms to all orders in $\alpha Z$, including logarithmically-enhanced terms that start at $\mathcal{O}(\alpha^{2}Z^{2}\log\epsilon_{\not{\pi}})$, as well as terms at $\mathcal{O}(\alpha^{2}Z\log\epsilon_{\not{\pi}})$, and $\mathcal{O}(\alpha^{2}\log\epsilon_{\not{\pi}})$ can be captured, see Refs. Jaus and Rasche (1970); Jaus (1972); Jaus and Rasche (1987); Sirlin (1978); Sirlin and Zucchini (1986); Sirlin (1987) and, in an EFT formalism, Refs. Ando et al. (2004); Cirigliano et al. (2023b); Hill and Plestid (2024a, b); Borah et al. (2024). Subleading interactions, such as the interactions of the photon with the nucleon magnetic moment, are proportional to the ultrasoft momentum and appear at $\mathcal{O}(\alpha\epsilon_{\text{recoil}})$ beyond the order at which we work.
Potential modes: Through topology 1$(c)$, potential modes give rise to $\mathcal{O}(\alpha\epsilon_{\not{\pi}})$ and $\mathcal{O}(\alpha\epsilon_{\chi})$ corrections to $\delta_{\text{NS}}$. The former depend on the electron energy ($E_{e}$) and are induced by diagrams with the EW vector current and the EM charge density. The latter are $E_{e}$-independent and are induced by the axial current and the nucleon magnetic moments or recoil corrections to the vector current. Three-body (3b) potentials contribute at $\mathcal{O}(\alpha\epsilon_{\chi}^{2})$ and are not shown.
@@
Hard modes: Hard modes give $\mathcal{O}(\alpha)$ corrections to $g_{V}$ Cirigliano et al. (2023b) and generate the $\mathcal{O}(\alpha\epsilon_{\chi})$ two-nucleon counterterms ($g^{N\!N}_{V1,V2}$) needed for renormalization. In addition, they produce $\mathcal{O}(\alpha\epsilon_{\not{\pi}},\alpha\epsilon_{\chi})$ effects in $\delta_{\text{NS}}$ through the electromagnetic pion mass splitting in pion-mediated 2b currents, see diagram 1$(d)$. The pion-mass splitting corrections are the nuclear analogs of the pion-induced RC in neutron decay Cirigliano et al. (2022b).
The implication of this analysis is that in chiral EFT the dominant contribution to $\delta_{\text{NS}}$ comes from the matrix element of appropriate EW potentials between the initial and final nuclear states. Some contributions (from pion exchange) do not arise from nuclear $\gamma W$ box diagrams. Sensitivity of $\delta_{\text{NS}}$ to intermediate nuclear states, including low-lying levels, arises from ultrasoft contributions that start to $\mathcal{O}(\alpha\epsilon_{\text{recoil}})$. This result is seemingly at odds with a recent dispersive analysis Seng and Gorchtein (2023a); Gorchtein and Seng (2024) in which some individual contributions scale as $\mathcal{O}(\alpha\sqrt{\epsilon_{\text{recoil}}})$, and thus enhanced compared to the identified EFT scalings. We therefore turn next to a detailed comparison to the dispersive representation.
@@
One way to perform the loop integral in Eq. (Radiative corrections to superallowed $\boldsymbol{\beta}$ decays in effective field theory) relies on a Wick rotation $\nu\to i\nu_{E}$, which is advantageous when the Compton tensor is expressed via a dispersion relation. However, as pointed out in Ref. Gorchtein and Seng (2024), such a Wick rotation is not always possible. In the case of low-lying nuclear states, as does happen in the ${}^{10}\text{C}\to{}^{10}\text{B}$ decay, the additional pole can move into the third quadrant and thus must be subtracted explicitly. It was found that the resulting residue contribution becomes singular for $E_{e}\to 0$, which could lead to a numerical enhancement. Such an enhancement should be reflected by the momentum scaling and a different region in the EFT analysis.
@@
where $s=M^{2}+\nu^{2}-\mathbf{q}^{2}+2M\nu$ and $M^{2}-\bar{M}^{2}=2M\Delta$. Here, $g_{A}$ and $g_{M}$ parameterize the matrix elements for the interaction with the EW and EM current, respectively. We focus on a single intermediate state with mass $\bar{M}$, with $\Delta>0$ corresponding to a low-lying state. The prefactor has been chosen to match the corresponding EFT expression Cirigliano et al. (2024b), counting the binding energy $\Delta\simeq q_{\text{ext}}$ as before.
We can evaluate the integral by collecting all three residues in the upper half plane, see Ref. Cirigliano et al. (2024b), which gives
@@
which is again finite for $E_{e}\to 0$, but, contrary to Eq. (12), scales as $\mathcal{O}(\alpha\sqrt{\epsilon_{\text{recoil}}})$ and could thus be enhanced numerically. The solution to this apparent mismatch is that the Wick-rotated integral also involves terms scaling with $\sqrt{\Delta}$, and one can show explicitly that Cirigliano et al. (2024b)
@@
Leading contributions to $\delta_{\emph{NS}}$.—In Ref. Cirigliano et al. (2024b) we derive the nuclear decay rate in the EFT framework, while here we focus on the implications for $\delta_{\text{NS}}$. Potential modes induce an effective Hamiltonian of the form
@@
with the explicit expressions for $\mathcal{V}_{E}$, $\mathcal{V}_{E}^{\pi}$, and $\mathcal{V}_{m_{e}}^{\pi}$ given in Ref. Cirigliano et al. (2024b). The latter two depend on the pion mass splitting, $M_{\pi^{\pm}}^{2}-M_{\pi^{0}}^{2}=2e^{2}F_{\pi}^{2}Z_{\pi}$, encoding effects of hard photons. These potentials are energy dependent and affect both the spectral shape and the total decay rate at $\mathcal{O}(\alpha\epsilon_{\not{\pi}})$.
Additional contributions arise from diagrams 1$(c,d)$ when using subleading vertices instead of inserting a lepton momentum. One order down in the chiral expansion at $\mathcal{O}(\alpha\epsilon_{\chi})$ we obtain potentials that are independent of the lepton momenta and thus contribute to $\mathcal{V}^{0}$. These $\mathcal{O}(\alpha\epsilon_{\chi})$ terms can be further decomposed as
@@
corresponding to diagrams 1$(c,d)$ via magnetic, recoil, and contact-term contributions Cirigliano et al. (2024b). Beyond tree level, the potentials (16) and (17) receive corrections from soft pions and photons at $\mathcal{O}(\alpha\epsilon_{\chi}^{2})$ and $\mathcal{O}(\alpha^{2})$, beyond the accuracy of this work. At this order we also expect effects from 3b potentials.
@@
$\mathcal{V}_{0}^{\text{CT}}$ depends on two unknown LECs and corresponds to genuine new 2b contributions arising from high-momentum photon exchange. It is an intrinsic two-nucleon effect that cannot be obtained from one-nucleon processes. Below we compute the contributions of $g^{N\!N}_{V1,V2}$ by using the scaling discussed below Eq. (6) for the LECs and treating the result as an overall uncertainty Cirigliano et al. (2024b).
In the EFT approach, up to the order considered, $\delta_{\text{NS}}$ is entirely determined by matrix elements of appropriate potentials, see Eq. (15), between the initial and final states without dependence on intermediate nuclear states. The EFT power counting indicates that $\delta_{\text{NS}}$ receives a LO $E_{e}$-independent contribution of $\mathcal{O}(\alpha\epsilon_{\chi})$, $\delta_{\text{NS}}^{(0)}$, and an $E_{e}$-dependent contribution of $\mathcal{O}(\alpha\epsilon_{\not{\pi}})$, $\delta^{E}_{\text{NS}}$. In the case of $\delta_{\text{NS}}^{(0)}$ we also found an $\mathcal{O}(\alpha^{2})$ potential $\mathcal{V}_{+}$ that needs to be included for $\mathcal{O}(10^{-4})$ precision Cirigliano et al. (2024b).